Consider the following project:

Task ID Duration (Days) Predecessors
2 70 -
3 30 2
4 120 3
5 120 2
6 120 5
7 120 3,6
8 150 3,6
9 115 3,6
10 130 7,9
11 90 7,9
12 30 7,9
13 120 7,9
14 75 10
15 30 14
16 85 6
17 165 16
18 90 16
19 90 17,18
20 20 8,11,13,15,19

What is the slack for task 8?
Answer #1

Answer:

step 1 find early start, and early finish from the start of the project for each activity
step 2 Find Maximum EF and keep it as the LF of last activities. Max of EF's is 685
step 3 calculate LS for each activity after calculating its EF
Activity Duration, D Early start, ES=Max of early finish of preceding activities Early finish, EF = ES + D Late finish, LF= Min of LS of successor activities Late start, LS= LF - D Total slack= LF-EF
2 70 0 70
3 30 70 100
4 120 100 220
5 120 70 190
6 120 190 310
7 120 310 430
8 150 310 460 665 515 205
9 115 310 425
10 130 430 560
11 90 430 520
12 30 430 460
13 120 430 550
14 75 560 635
15 30 635 665
16 85 310 395
17 165 395 560
18 90 395 485
19 90 560 650
20 20 665 685 685 665 0

the slack for task 8 is 205